# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v10

@@ 10. Elementary school survey July 1919 2
A unique feature of the survey methodology is the use of young researchers as commissioned researchers. The household budget survey and leisure life survey were conducted by Yasunosuke Gonda, a faculty member of the Doitsugaku kyōkai gakkō (Association for German Sciences), who had researched cinematograph and entertainment. Tetsuo Hoshino, a medical doctor, was responsible for the sanitary survey. The labor survey was managed by Yoshitsuru Yamana, a graduate of Tokyo Imperial University who contributed to the organization of labor unions. In addition, Toyotaro Miyoshi, a sociology student at the Tokyo Imperial University, and others became temporary employees (see footnote 55). At the time, the Health Bureau of the Ministry of the Interior conducted several sanitary surveys of rural areas. Takano may have designed the urban sanitary survey from this background (Kawai 1981a, pp. 8–9). However, based on the idea that “health and sanitation surveys can at least encompass economic and social surveys,” Takano included a wide range of survey items, not limited to sanitation surveys (Department of Health, Ministry of the Interior 1923a, p. 56). Kawai (1980, p. 55) credits the Tsukishima Survey as a noteworthy pioneering concept: “going beyond a mere insurance and sanitation survey, it attempted to accurately depict the state of urban workers’ lives as they became established in the city during modernization, with skilled workers as its core.”
@@ B.6.1 Daily Wage Distribution by Ages
Next, I shift the predicted distribution based on the difference in the average daily wages in the ASCT and LST. The weighted average of the daily wage is $2.56$ yen for the LST, whereas that for the ASCT is $1.05$ yen, meaning that there is approximately $1.5$ yen difference. [^99] For the LST, I use the number of factory workers in each age bin as weight. The ASCT reports the average daily wage for four different categories of factories in each manufacturing sector: the factory scale (ten+ or below ten workers) and engines (use or not). I calculate the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights. Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.100100100This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar. In the LST (i.e., the statistic on which the predicted distribution is based), the weighted average for those aged $14$ and older was $3.25$ yen. In comparison, the overall weighted average is $3.23$ yen, a difference of only $0.02$ yen. A similar trend can be observed in the statistics obtained from the ASCT: the weighted average for those aged $15$ and older is $1.14$, while the overall average is $1.13$, a difference of only $0.01$ yen. This suggests that the wage distribution in the ASCT is similar to that in the LST. Figure B.8a illustrates this operation: dotted curve is a fractional polynomial shifted from the original fractional polynomial shown as a short dashed curve.
@@
The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12–14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial.101101101The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$. I used the 25–65 age range for the prediction based on the age bins used in the LST. Let this curve predict wage distribution for the true distribution. The average daily wage for the 30–40-year-old bin in the shifted fractional polynomial is $1.55$ yen, which is similar to that for the same-age bin in the predicted wage distribution (i.e., $1.52$ yen). However, the wage gaps for the younger ages show more significant differences. If our interest is in the late 20s to the early 40s, it is plausible to employ the predicted wages for the wage comparison in Section 3.
@@ B.6.3 Estimating Monthly Income of Adult Male Factory Workers
where $w^{\textit{Regular}}>0$ is the monthly regular wage and $w^{\textit{Ancillary}}\geq 0$ indicates the monthly ancillary wage. Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$.102102102Strictly, this weight may not be required to be in this range. I use this range because all the cases reported in Kitazawa (1924) take the weights between zero and one. The monthly wage equation is then specified as follows:
@@ B.7 Measurement Errors in the Unadjusted Panel Dataset
B.8 Distribution of the Idiosyncratic Shocks
@@
Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.103103103I use the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets. This is because the daily panels include a set of consecutive household-date cells with no income. These cells cannot be used to predict the idiosyncratic shocks as the residuals are economically meaningfully, and thus it is practically impossible to illustrate the idiosyncratic shocks based on the daily dataset. The idiosyncratic shocks are calculated as residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as follows:
@@
B.9 Unit Root and Autocorrelations
@@ B.10 Risk Preferences
Schulhofer-Wohl (2011) shows a specific case (utility function) that the heterogeneous risk preference may lead to bias in the estimation of the income elasticity under the standard risk-sharing regressions and proposes a parametric test for that heterogeneity. This test is valid under a particular utility function. Specifically, the bias could be raised in the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock cannot be removed using the time-fixed effect in the risk-sharing regressions.104104104Specifically, Schulhofer-Wohl (2011, pp. 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$. The log-transformed first-order condition becomes: $\text{log}c_{i,t}=\frac{\text{log}\dot{\alpha}_{i}}{\dot{\gamma}_{i}}-\frac{\text{log}\dot{\lambda}_{t}}{\dot{\gamma}_{i}},$ (18) where $\dot{\lambda}_{t}$ indicates the Lagrangian multiplier. This result implies that the effects of the aggregate shocks ($\dot{\lambda}_{t}$) depend on the household’s attitudes toward risk and that the first-differencing does not remove the second term in equation 18. As explained, this setting differs from the one used in this current study. Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences. Their test requires a panel dataset with a very long time dimension that is not applicable to my THBS dataset.105105105The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985. The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p. 453). See also Shrinivas and Fafchamps (2018) and Mozzoco and Saini (2018) for the subsequent discussions.
An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers. Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.106106106This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there was a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P. 931). They found that the high-risk-tolerance workers have more procyclical incomes than the low-risk-tolerance workers.
@@
$d_{i}$ is an indicator variable that takes one if a household $i$ belongs to a specific group that might have had different preferences on their jobs than the other heads in the sample. Firstly, I test the hypothesis that the heads who work in the large-scale factories have different risk preferences. Although there are no specific anecdotes behind this, the workers in the large-scale factories might have had a lower risk tolerance than those in the small- and medium-scale enterprises. Thirteen heads who worked in the Ishikawajima shipyard and governmental factory are classified into this group. Secondly, I test the hypothesis that the heads who work in the smithing and the non-machinery factories have different risk preferences. As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.107107107Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6). However, my THBS sample does not include the workers in canning factories. The workers in non-machinery factories might also have different job preferences than the machinery factory workers. This group has only three heads: one worked in a smithing factory, and two worked in food and paper-making factories.
@@ C.1 Conceptual Framework
where $\tilde{y}_{i,t}$ is the head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term. I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.108108108In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2). The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue arises if I use the disposable income in this equation. The regression is separately run for each component in the net income vector ($\bm{z}$).

### v11

@@ 10. Elementary school survey July 1919 2
A unique feature of the survey methodology is the use of young researchers as commissioned researchers. The household budget survey and leisure life survey were conducted by Yasunosuke Gonda, a faculty member of the Doitsugaku kyōkai gakkō (Association for German Sciences), who had researched cinematograph and entertainment. Tetsuo Hoshino, a medical doctor, was responsible for the sanitary survey. The labor survey was managed by Yoshitsuru Yamana, a graduate of Tokyo Imperial University who contributed to the organization of labor unions. In addition, Toyotaro Miyoshi, a sociology student at the Tokyo Imperial University, and others became temporary employees (see footnote 54). At the time, the Health Bureau of the Ministry of the Interior conducted several sanitary surveys of rural areas. Takano may have designed the urban sanitary survey from this background (Kawai 1981a, pp. 8–9). However, based on the idea that “health and sanitation surveys can at least encompass economic and social surveys,” Takano included a wide range of survey items, not limited to sanitation surveys (Department of Health, Ministry of the Interior 1923a, p. 56). Kawai (1980, p. 55) credits the Tsukishima Survey as a noteworthy pioneering concept: “going beyond a mere insurance and sanitation survey, it attempted to accurately depict the state of urban workers’ lives as they became established in the city during modernization, with skilled workers as its core.”
@@ B.6.1 Daily Wage Distribution by Ages
Next, I shift the predicted distribution based on the difference in the average daily wages in the ASCT and LST. The weighted average of the daily wage is $2.56$ yen for the LST, whereas that for the ASCT is $1.05$ yen, meaning that there is approximately $1.5$ yen difference. [^96] For the LST, I use the number of factory workers in each age bin as weight. The ASCT reports the average daily wage for four different categories of factories in each manufacturing sector: the factory scale (ten+ or below ten workers) and engines (use or not). I calculate the weighted average of the daily wage using the number of factory workers in each age bin (12–14 and 15+) and in each category as weights. Thus, I shift the functional polynomial function by this amount of difference to yield the proxy distribution for 1918.979797This operation, therefore, assumes that the wage-by-age distributions in 1918 and 1924 are reasonably similar. In the LST (i.e., the statistic on which the predicted distribution is based), the weighted average for those aged $14$ and older was $3.25$ yen. In comparison, the overall weighted average is $3.23$ yen, a difference of only $0.02$ yen. A similar trend can be observed in the statistics obtained from the ASCT: the weighted average for those aged $15$ and older is $1.14$, while the overall average is $1.13$, a difference of only $0.01$ yen. This suggests that the wage distribution in the ASCT is similar to that in the LST. Figure B.8a illustrates this operation: dotted curve is a fractional polynomial shifted from the original fractional polynomial shown as a short dashed curve.
@@
The long dashed curve in Figure B.8a examines this relationship by fitting another fractional polynomial model using the 12–14-year old bin figure of the ASCT and the predicted wages for 25–65 from the shifted fractional polynomial. [^98] The fitted model is as follows: $\hat{y}^{\textit{Daily Wage}}=-1.19+0.14a^{1}-0.002a^{2},\quad R^{2}=0.98.$ The deviance is reported to be $-149$. I used the 25–65 age range for the prediction based on the age bins used in the LST. Let this curve predict wage distribution for the true distribution. The average daily wage for the 30–40-year-old bin in the shifted fractional polynomial is $1.55$ yen, which is similar to that for the same-age bin in the predicted wage distribution (i.e., $1.52$ yen). However, the wage gaps for the younger ages show more significant differences. If our interest is in the late 20s to the early 40s, it is plausible to employ the predicted wages for the wage comparison in Section 3.
@@ B.6.3 Estimating Monthly Income of Adult Male Factory Workers
where $w^{\textit{Regular}}>0$ is the monthly regular wage and $w^{\textit{Ancillary}}\geq 0$ indicates the monthly ancillary wage. Online Appendix B.6.2 shows that $w^{\textit{Ancillary}}$ can be calculated as the product of the monthly regular wage and weight for the ancillary wage, $\vartheta\in(0,1)$. [^99] Strictly, this weight may not be required to be in this range. I use this range because all the cases reported in Kitazawa (1924) take the weights between zero and one. The monthly wage equation is then specified as follows:
@@ B.7 Measurement Errors in the Unadjusted Panel Dataset
B.8 Testing Attenuation Bias in the Unadjusted Panel Dataset
In this subsection, I assess the validity of the aggregation method used in this study. Adjusting the aggregate data to account for the timing of paydays reveals more transparent relationships between income and expenditure than the unadjusted data (Section 3.3). To demonstrate this mechanism, I listed the estimates from the unadjusted monthly panel dataset in Column (2) of Table B.5. The estimate for total consumption is $0.26$, which is much lower than the estimate for the adjusted series ($0.43$) in Column (1). In addition, most subcategories show systematically smaller estimates than those from the adjusted data listed in Column (2). For example, the estimate for the food category is now $0.10$ and statistically insignificant. This means that measurement errors in the unadjusted aggregation lead to strong attenuation bias, which results in false negatives (i.e., false rejections of the null of full risk-sharing hypothesis) in most of the consumption categories. I, therefore, specified a generating process of systematic measurement errors in the aggregate survey data, which has attracted attention in the macroeconomics literature (Cochrane 1991; Dynarski et al. 1997; Gervais and Klein 2010; Nelson 1994). Moreover, several subcategories select values close to one or large negative values, which are no longer economically interpretable.100100100For example, the estimate for housing was negative. The households paid a fixed rent amount monthly. Given that the rent payday was fixed, the shifts in wage paydays might create cells with less income with a fixed amount of rent. In the case of seasonal consumption categories, the shifts may coincide with the timing of seasonal spending. The THBS households tended to purchase work clothes in March to prepare for the start of the new fiscal year in April. They also spent more on clothing in June, likely to buy summer clothes. The shift of paydays to these months may have created a strong but false positive correlation. This means that the miss-assignment of paydays in the unadjusted data may cause critical statistical inference issues in a few subcategories.
Table B.5: Testing Attenuation Bias in the Unadjusted Panel Dataset
(1) Adjusted Monthly Panels (2) Unadjusted Monthly Panels
Disposable income Obs. Disposable income Obs.
Coef. Std. error Coef. Std. error
Total consumption 0.433 [0.059]*** 123 0.261 [0.054]*** 141
      Food 0.293 [0.066]*** 114 0.100 [0.072] 132
      Housing 0.013 [0.084] 105 -0.263 [0.167] 124
      Utilities 0.343 [0.236] 110 0.197 [0.334] 130
      Furniture -0.114 [0.706] 98 0.250 [0.339] 112
      Clothes 0.347 [0.424] 112 0.806 [0.262]*** 132
      Education 0.162 [0.128] 110 0.069 [0.146] 128
      Medical expenses 0.194 [0.120] 114 -0.128 [0.094] 132
      Entertainment expenses 0.716 [0.181]*** 113 0.194 [0.153] 132
      Transportation 0.605 [0.264]** 97 0.362 [0.291] 115
      Gifts 0.469 [0.197]** 113 0.622 [0.306]* 130
      Miscellaneous 0.522 [0.252]** 110 0.246 [0.267] 126
Notes: This table shows the results of Equation 1. The regressions of the 11 measures of log-transformed consumption on log-transformed disposable income, the family size control, household-fixed effect, and time-fixed effect are listed. Columns 1 and 2 show the results for the adjusted monthly and unadjusted monthly panel datasets, respectively. The estimates in column (1) are the same as those reported in Table 4. The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
B.9 Considering Child Share as a Preference Shock Variable
As explained in the main text, the child share variable exhibits significantly more within-variations than that of the family size. Table B.6 presents the results from specifications that include the share of children under the primary school age range, specifically those aged 0–5, as an alternative preference shock variable. Panels A and B show the results for the consumption smoothing and risk-coping strategy regressions, respectively. The results are materially similar to the main results presented in the main text (Tables 4 and 5). Note that the results remain robust to different definitions of age bins (not reported).
Table B.6: Results for the Alternative Specification:
Conditioning on the Share of Children Aged 0–5
(1) Semi-monthly Panels (2) Adjusted Monthly Panels
Disposable income Obs. Disposable income Obs.
Panel A: Consumption Smoothing Coef. Std. error Coef. Std. error
Total consumption 0.361 [0.034]*** 278 0.440 [0.058]*** 123
      Food 0.130 [0.028]*** 259 0.277 [0.062]*** 114
      Housing 0.097 [0.413] 134 0.015 [0.079] 105
      Utilities -0.005 [0.140] 223 0.339 [0.232] 110
      Furniture 0.431 [0.234]* 169 0.172 [0.665] 98
      Clothes 0.381 [0.189]* 238 0.319 [0.412] 112
      Education 0.103 [0.050]** 236 0.166 [0.127] 110
      Medical expenses 0.037 [0.066] 258 0.182 [0.122] 114
      Entertainment expenses 0.304 [0.085]*** 250 0.695 [0.176]*** 113
      Transportation 0.281 [0.110]** 187 0.613 [0.267]** 97
      Gifts 0.524 [0.107]*** 234 0.476 [0.192]** 113
      Miscellaneous -0.079 [0.148] 214 0.506 [0.253]* 110
(1) Semi-monthly Panels (2) Adj. Monthly Panels
Head’s earning Obs. Head’s earning Obs.
Panel B: Risk-Coping Strategy Coef. Std. error Coef. Std. error
Panel B-1: Savings, insurance, borrowing, and gifts (yen)
Net savings -0.043 [0.018]** 289 -0.035 [0.028] 124
Net insurance 0.001 [0.013] 289 0.002 [0.032] 124
Net borrowing -0.021 [0.027] 289 -0.026 [0.017] 124
Net credit purchase -0.129 [0.039]*** 289 -0.021 [0.042] 124
Net gifts -0.052 [0.020]** 289 -0.006 [0.029] 124
Panel B-2: Labor supply adjustments & sales of miscellaneous assets (yen)
Other member’s earnings 0.001 [0.010] 289 0.036 [0.033] 124
Sales of miscellaneous assets -0.000 [0.000] 289 0.000 [0.001] 124
Notes: Panel A: This table shows the results of Equation 1. The regressions of the 11 measures of log-transformed consumption on log-transformed disposable income, the proportion of children aged 0–5, household-fixed effect, and time-fixed effect are listed. Columns 1 and 2 show the results for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
Panel B: This table shows the results of equation 2. The proportion of children aged 0–5, household fixed effects, and time-fixed effects are included in all the regressions. Columns 1 and 2 show the estimates for the semi-monthly and adjusted monthly panel datasets, respectively. The estimated coefficients on the head’s earnings are listed in the “Coef.” columns.
Panel B-1: Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts.
Panel B-2: “Other member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of miscellaneous assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles. All dependent variables are in yen.
B.10 Distribution of the Idiosyncratic Shocks
@@
Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset.101101101I use the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets. This is because the daily panels include a set of consecutive household-date cells with no income. These cells cannot be used to predict the idiosyncratic shocks as the residuals are economically meaningfully, and thus it is practically impossible to illustrate the idiosyncratic shocks based on the daily dataset. The idiosyncratic shocks are calculated as residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as follows:
@@
B.11 Unit Root and Autocorrelations
@@ B.12 Risk Preferences
Schulhofer-Wohl (2011) shows a specific case (utility function) that the heterogeneous risk preference may lead to bias in the estimation of the income elasticity under the standard risk-sharing regressions and proposes a parametric test for that heterogeneity. This test is valid under a particular utility function. Specifically, the bias could be raised in the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock cannot be removed using the time-fixed effect in the risk-sharing regressions.102102102Specifically, Schulhofer-Wohl (2011, pp. 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$. The log-transformed first-order condition becomes: $\text{log}c_{i,t}=\frac{\text{log}\dot{\alpha}_{i}}{\dot{\gamma}_{i}}-\frac{\text{log}\dot{\lambda}_{t}}{\dot{\gamma}_{i}},$ (18) where $\dot{\lambda}_{t}$ indicates the Lagrangian multiplier. This result implies that the effects of the aggregate shocks ($\dot{\lambda}_{t}$) depend on the household’s attitudes toward risk and that the first-differencing does not remove the second term in equation 18. As explained, this setting differs from the one used in this current study. Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences. Their test requires a panel dataset with a very long time dimension that is not applicable to my THBS dataset.103103103The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985. The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p. 453). See also Shrinivas and Fafchamps (2018) and Mozzoco and Saini (2018) for the subsequent discussions.
An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers. Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.104104104This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there was a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P. 931). They found that the high-risk-tolerance workers have more procyclical incomes than the low-risk-tolerance workers.
@@
$d_{i}$ is an indicator variable that takes one if a household $i$ belongs to a specific group that might have had different preferences on their jobs than the other heads in the sample. Firstly, I test the hypothesis that the heads who work in the large-scale factories have different risk preferences. Although there are no specific anecdotes behind this, the workers in the large-scale factories might have had a lower risk tolerance than those in the small- and medium-scale enterprises. Thirteen heads who worked in the Ishikawajima shipyard and governmental factory are classified into this group. Secondly, I test the hypothesis that the heads who work in the smithing and the non-machinery factories have different risk preferences. As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.105105105Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6). However, my THBS sample does not include the workers in canning factories. The workers in non-machinery factories might also have different job preferences than the machinery factory workers. This group has only three heads: one worked in a smithing factory, and two worked in food and paper-making factories.
@@ C.1 Conceptual Framework
where $\tilde{y}_{i,t}$ is the head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term. I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.106106106In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2). The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue arises if I use the disposable income in this equation. The regression is separately run for each component in the net income vector ($\bm{z}$).